# Covariate Adjusted Response Adaptive Design with Delayed Outcomes

### v1

@@ Abstract
Covariate-adjusted response adaptive (CARA) designs have gained widespread adoption for their clear benefits in enhancing experimental efficiency and participant welfare. These designs dynamically adjust treatment allocations during interim analyses based on participant responses and covariates collected during the experiment. However, delayed responses can significantly compromise the effectiveness of CARA designs, as they hinder timely adjustments to treatment assignments when certain participant outcomes are not immediately observed. In this manuscript, we propose a fully forward-looking CARA design that dynamically updates treatment assignments throughout the experiment as response delay mechanisms are progressively estimated. Our design strategy is informed by novel semiparametric efficiency calculations that explicitly account for outcome delays in a multi-stage adaptive experiment. Through both theoretical investigations and simulation studies, we demonstrate that our proposed design offers a robust solution for handling delayed outcomes in CARA designs, yielding significant improvements in both statistical power and participant welfare.
Keywords: Causal inference; Delayed outcomes; Frequentist adaptive experimental design; Response adaptive designs
@@ 1.1 Motivation and our contribution
In clinical trials, field experiments in social sciences, and A/B tests, participants often enroll sequentially, and responses to treatments can vary due to individual-specific characteristics. In these contexts, covariate-adjusted response-adaptive (CARA) designs have gained increasing attention for their ability to enhance both the efficiency and the ethical or welfare considerations of experimental outcomes (Zhang et al., 2007; Hu et al., 2015; Robertson et al., 2023; Wei et al., 2023, 2024; Zhao, 2023). By dynamically adjusting treatment assignments based on accumulated responses and covariates during the experiment, CARA designs can be adopted to optimize treatment allocation to improve statistical power, reduce participant exposure to less effective treatments, enhance overall welfare, and maintain statistical validity for analyzing experimental data.
@@
Nevertheless, because CARA designs rely on participant responses collected during the experiment to adjust treatment assignments, delays in observing these responses can compromise the design’s effectiveness. This can be intuitively understood as follows: during interim analyses, if the primary outcomes for some participants are not yet observed, the experimental designer may lack the necessary information to optimize treatment assignments for future participants who have profiles similar to those with delayed responses. Furthermore, as is often the case in randomized experiments when participants are sequentially enrolled, the response delay mechanisms can vary among participants depending on their assigned treatment arm and baseline covariates. Below, we provide a motivating example with arm- and covariate-dependent response delays in a randomized experiment:
Example 1 (Arm- and covariate-dependent randomized experiments for a cash incentive trial among people living with HIV (PLHIV)).
In a randomized experiment conducted in Shinyanga region, Tanzania (Fahey et al., 2020), the goal is to assess the impact of cash incentives on improving retention in care and thus achieving viral suppression for PLHIV. In this experiment, 530 PLHIV aged 18 years or older were sequentially enrolled and randomly assigned to either receive cash incentives (treatment arm) or no cash incentives (control arm) between April 24, 2018, and December 14, 2018. Retention in care is evaluated by individual adherence to antiretroviral therapy measured by viral load six months after treatment assignment. However, delays in measuring viral load are observed, as the lab test requires clinical visits for blood draws, and some PLHIV miss their scheduled appointments. To investigate these response delay mechanisms, we record the date at which each patient completed six months of treatment and calculate the number of days delayed as the difference between the date of the viral load test and the six-month completion date. We further include biological sex as a baseline covariate to examine if the delay mechanism varies across females and males.
Figure 1: Distribution of the number of delayed days in viral load response collection across treatment arms and baseline covariate (i.e., biological sex) after completing six months of antiretroviral therapy.
Figure 1 presents a histogram showing the number of delayed days for viral load measurements, stratified by treatment and control arms as well as biological sex. The figure illustrates that the distribution of delayed days is influenced by both treatment assignment and biological sex. Specifically, the treatment arm shows a higher percentage of responses collected within 100 days compared to the control arm. Moreover, within the treatment arm, a larger proportion of female participants have response delays of fewer than 100 days compared to males. This observation suggests that for females receiving cash transfers, most viral load responses were collected within three months of completing the treatment. Overall, Figure 1 highlights that response delays in this randomized experiment are dependent on both the treatment arm and covariate.
To our knowledge, although there has been discussion about how response delays might affect statistical inference at the conclusion of CARA(Bai et al., 2008; Hu et al., 2008; Robertson et al., 2023), no previous CARA design has leveraged the delay mechanism to explicitly optimize experimental objectives. In this manuscript, we address the challenge of response delays in CARA and seek to answer the following questions: Does the optimal treatment allocation proposed in existing CARA literature remain effective for achieving design objectives? If not, what is the optimal treatment allocation? How can CARA be adapted to incorporate response delays for more effective attainment of these objectives? Finally, at the conclusion of the experiment, how can we construct an efficient statistical estimator that accounts for these delays? The following paragraphs outline our contributions in addressing these questions within the context of existing CARA literature.
@@
First, we demonstrate that the classical optimal allocation strategy used in existing CARA designs no longer provides optimal treatment assignment when accounting for arm- and covariate-dependent delays. Specifically, we first derive a semiparametric efficiency bound for treatment effect estimation accommodating both multi-valued treatments and outcome delay, which differ substantially from existing results in the literature (Hahn, 1998; Hirano et al., 2003; Cattaneo, 2010; Zhu and Zhu, 2023) (see Newey 1990; Van der Vaart 2000; Tsiatis 2006 for general discussions on semiparametric efficiency). By deriving semiparametric efficiency bounds that incorporate response delays in multi-stage CARA designs (Theorem 1), we show that the optimal delay mechanism involves a unique trajectory of treatment assignment probabilities that vary across different stages, resulting in treatment allocations that reduce variance for power maximization and minimize failure rates in welfare improvement designs (Figure 2). To our knowledge, this is the first design to explicitly explore delay mechanisms to optimize trial objectives, while existing literature focuses on conducting sequential tests (Hampson and Jennison, 2013), designing Bayesian adaptive designs where delay is independent of the arm and covariates (Lin et al., 2020), or conducting statistical inference after experiments with delayed responses have concluded (Bai et al., 2008; Hu et al., 2008).
@@
Second, because the delay mechanism is unknown a priori and only partially learnable during the experiment, we propose a fully forward-looking CARA design that sequentially updates treatment assignment probabilities to optimize the trial outcome by the experiment’s end while also estimating future delays to incorporate the experimental designer’s preferences. Our design encounters two design objectives frequently adopted in current CARA literature: maximizing experiment power (Tymofyeyev et al., 2007; Zhao, 2023) and improving overall participant welfare (or failure reduction) while maintaining statistical power constraints (Rosenberger et al., 2001; Robertson et al., 2023). We justify the benefits of our proposed forward-looking CARA design through theoretical investigation (Section 4) and a synthetic case study (Section 5). Under the design objective of maximizing experiment power, our proposed forward-looking CARA design attains higher estimation efficiency compared to the Neyman allocation and complete randomization designs in the presence of delayed responses. The efficiency gain is further justified by Theorem 5. Under the design objective of improving overall participant welfare, we demonstrate that our proposed forward-looking CARA design achieves a higher failure reduction rate, leading to greater welfare improvement compared to both the ethical design and the complete randomization design while maintaining statistical power. For both design objectives, our proposed design can provide valid statistical inference, which is justified in Theorem 2 and Theorem 3 and further demonstrated in a synthetic case study (Section 5).
@@
Third, we make theoretical contributions to the literature on CARA and adaptive experimental designs by establishing general conditions under which estimated treatment effects are consistent and asymptotically normally distributed (Lemma 1 and Theorem 2). Specifically, we do not restrict the potential outcomes or the delay mechanism to any parametric family of distributions, thereby alleviating the burden of choosing a particular set of parametric assumptions. Such flexibility is crucial for many applications, as the delay mechanism is typically arm- and covariate-dependent, as our earlier examples have demonstrated. Our theoretical development builds on martingale methods (Hall and Heyde, 1980) and a high-level condition (Condition 1) that the optimized allocations converge in large samples—an assumption satisfied by many adaptive experimental designs. As a result, the general consistency and asymptotic normality findings may be of independent interest. We then specialize in our proposed design and verify the design consistency condition (Section 4.3), providing a suite of tools for designing and analyzing adaptive experiments with delayed outcomes, along with rigorous statistical guarantees.
1.2 Setup: CARA with arm and covariate dependent delayed outcomes
@@
In the remainder of this introduction, we lay out the setup of CARA when responses are subject to delay. Formal statements of our assumptions are collected in Section 4. The experiment is conducted over $T$ stages, labeled $t=1,2,\dots,T$. In each stage $t$, $n_{t}$ participants are enrolled. Participants are indexed by $i=1,2,\dots,N$, with $N=\sum_{t=1}^{T}n_{t}$ being the total sample size. Occasionally, we will use $N_{t}$ to denote the cumulative sample size at the end of stage $t$, that is, $N_{t}=\sum_{s=1}^{t}n_{s}$. Therefore, $N=N_{T}$, where we omit the subscript to save notation. Finally, $S_{i}=t$ denotes that individual $i$ is enrolled in stage $t$.
@@
Upon enrollment, baseline covariates $X_{i}\in\mathcal{X}$ are collected for each participant, and they are then randomized into one of the treatment arms, denoted by $A_{i}\in\mathcal{A}$, which represents their actual treatment status. Following the Neyman-Rubin causal model, the potential outcomes are denoted by $Y_{i}(a)$ for different treatment $a\in\mathcal{A}$. In line with existing literature in adaptive design, we are in a scenario that the potential outcomes and the covariates, $(Y_{i}(a),X_{i})$, are independently and identically distributed across different stages. For future reference, we also introduce notation for mean potential outcomes: $\mu(x,a)=\mathbb{E}[Y_{i}(a)|X_{i}=x]$ and $\mu(a)=\mathbb{E}[Y_{i}(a)]$, [^1] Other moments of the potential outcomes are $\mu_{s}(x,a)=\mathbb{E}[Y_{i}(a)^{s}|X_{i}=x]$ and $\mu_{s}(a)=\mathbb{E}[Y_{i}(a)^{s}]$, but we omit the subscript whenever $s=1$ for ease of notation. and the conditional and unconditional treatment effects are defined as
$\tau(x,a^{\prime},a)$ $=\mu(x,a^{\prime})-\mu(x,a)\quad\text{and}\quad\tau(a^{\prime},a)=\mu(a^{\prime})-\mu(a).$
@@
In this manuscript, we allow the delay mechanism $D_{i}$ to depend on the covariates and the treatment assignment status, and is independent of the potential outcomes conditional on the covariates $X_{i}$ and the treatment assignment $A_{i}$. Similar to the assumption of identical distribution of potential outcomes across stages mentioned earlier, we assume that the delay distribution does not shift over time. Together, these can be formally expressed as:
@@
As the delay distribution does not shift over time, it allows us to sequentially learn the delay mechanism using accrued data and to optimize treatment allocation. Finally, we note that this assumption does not preclude the possibility that the delay mechanism may depend on the observed outcomes.
In CARA designs, the treatment assignment is sequentially updated based on accumulated data to achieve a pre-specified design objective. Specifically, at stage $t$, suppose we have collected historical participant information, denoted as
$\mathcal{H}_{t}=\{(A_{i},X_{i},Y_{i},D_{i})\}_{i=1}^{N_{t}},$
where we recall that the sample indexed by $i=1,2,\dots,N_{t}$ consists of all participants enrolled in the first $t$ stages. The treatment assignment for the next stage, $t+1$, denoted as $\hat{e}^{*}_{t+1}(a|x)$, is determined by both this historical information and the covariate information of the newly enrolled participants in Stage $t+1$. Formally, this is captured by
$\mathbb{P}[A_{i}=a|X_{i}=x,S_{i}=t+1,\mathcal{H}_{t}]=\hat{e}^{*}_{t+1}(a|x),$
which means the treatment assignment adapts to $\mathcal{H}_{t}$.
@@ 2 Optimal delay-adjusted treatment allocation in CARA
In this section, we discuss how delayed response in CARA affects optimal treatment allocation rules. Compared to the existing literature on CARA or response adaptive design (RAR) with delayed outcomes (Bai et al., 2008; Hu et al., 2008), to our knowledge, no prior designs have explicitly leveraged delay mechanisms to optimize experimental objectives.
@@
While the design goals of CARA vary across applications, maintaining statistical power to assess the treatment’s effectiveness on the primary outcome remains a central concern. A statistically efficient estimator of the treatment effect also contributes to a robust adaptive design, as the estimated treatment effects are used to sequentially revise the treatment allocation strategy. However, it remains unclear what constitutes a good estimator of the treatment effect in the presence of delayed outcomes. Moreover, the semiparametric efficiency bound for estimating the treatment effect in multi-stage experiments with delayed outcomes has not been developed. To provide practical guidance for designing response-adaptive experiments, we first present a new result on the semiparametric efficiency bound of estimating the average treatment effect, defined as the mean difference between the potential outcomes of two treatment arms $a$ and $a^{\prime}$. In such experiments with delayed outcomes, we have
$\int\limits_{x\in\mathcal{X}}p(x)\Bigg{[}\frac{\sigma^{2}(x,a^{\prime})}{\sum\limits_{t=1}^{T}r_{t}e_{t}(a^{\prime}|x)\rho(T-t|x,a^{\prime})}$ $+\frac{\sigma^{2}(x,a)}{\sum\limits_{t=1}^{T}r_{t}e_{t}(a|x)\rho(T-t|x,a)}+\big{(}\tau(x,a^{\prime},a)-\tau(a^{\prime},a)\big{)}^{2}\Bigg{]}\mbox{d}x,$ (1)
@@
where $r_{t}=\mathbb{P}[S_{i}=t]$, the fraction of participants enrolled in stage $t$, $p(x)$ is the probability density function of the covariate $X$, and $\sigma_{t}^{2}(x,a)$ is the conditional variance of potential outcomes under treatment $a\in\{0,1\}$ at time $t$. See Section 4, Theorem 1 for a rigorous statement.
@@
From the above result, it is evident that when the delay mechanism of the response depends on the covariates and treatment status, the optimal treatment allocation rules in classical CARA designs must be modified to account for the impact of the delay mechanism. In what follows, we examine how the optimal treatment allocation rule is affected by the arm- and/or covariate-dependent delayed responses in two common experimental objectives: one aimed at enhancing statistical power (Tymofyeyev et al., 2007), and the other focused on minimizing the expected number of failures while maintaining a minimum power requirement for binary outcomes (Rosenberger et al., 2001). In both design goals, the variance of the estimated treatment effects plays an essential role. To simplify the presentation, we focus on two-arm CARA designs with $a\in\{0,1\}$. For multi-arm experiments, the design objective can be generalized using noncentrality parameters, as discussed in Lachin (1977); Tymofyeyev et al. (2007).
@@
Design objective 1: Optimal treatment allocation in power maximization design with delayed response. This design aims to sequentially assign treatments to minimize the variance of the estimated average treatment effect (ATE) for two treatment arms, denoted by $a\in\{0,1\}$, thereby maximizing statistical power. The optimal allocation is determined by solving the following optimization problem:
@@
$\min_{e_{1},\ldots,e_{T}}\ \mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))\quad\text{subject to}\quad\delta\leq e_{t}(\cdot)\leq 1-\delta\quad\text{for }t=1,\ldots,T,$ (2)
@@
where $\delta$ ensures that the assignment probabilities are bounded away from 0 and 1, maintaining sufficient randomness in treatment assignments. The objective function $\mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))$ is a component of the semiparametric efficiency bound related to the propensity scores, and is defined as:
@@
$\mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))=\sum_{x\in\mathcal{X}}p(x)\left[\frac{\sigma_{t}^{2}(x,1)}{\sum_{t=1}^{T}r_{t}e_{t}(1|x)\rho(T-t|x,1)}+\frac{\sigma_{t}^{2}(x,0)}{\sum_{t=1}^{T}r_{t}e_{t}(0|x)\rho(T-t|x,0)}\right].$

### v2

# Xinwei Ma Jingshen Wang Waverly Wei

@@ Abstract
Covariate-adjusted response-adaptive (CARA) designs have gained widespread adoption for their clear benefits in enhancing experimental efficiency and participant welfare. These designs dynamically adjust treatment allocations during interim analyses based on participant responses and covariates collected during the experiment. However, delayed responses can significantly compromise the effectiveness of CARA designs, as they hinder timely adjustments to treatment assignments when certain participant outcomes are not immediately observed. In this paper, we propose a fully forward-looking CARA design that dynamically updates treatment assignments throughout the experiment as response delay mechanisms are progressively estimated. Our design strategy is informed by novel semiparametric efficiency calculations that explicitly account for outcome delays in a multi-stage setting. Through both theoretical investigations and simulation studies, we demonstrate that our proposed design offers a robust solution for handling delayed outcomes in CARA designs, yielding significant improvements in both statistical power and participant welfare.
Keywords: Delayed outcomes; Frequentist adaptive experimental design; Response adaptive designs.
@@ 1.1 Motivation and contribution
In clinical trials, social science field experiments, and A/B tests, participants often enroll sequentially, and responses to treatments can vary due to individual-specific characteristics. In these contexts, covariate-adjusted response-adaptive (CARA) designs have gained increasing popularity: by dynamically adjusting treatment assignments based on accumulated covariates and outcome information during the experiment, CARA designs can be adopted to optimize treatment allocation to improve statistical power (Hahn, Hirano and Karlan, 2011; Hu, Zhu and Hu, 2015; Blackwell, Pashley and Valentino, 2023), reduce participant exposure to less effective treatments and enhance overall welfare (Hu and Rosenberger, 2006; Hu, Zhu and Hu, 2015; Wei, Ma and Wang, 2024, 2025), and maintain statistical validity for analyzing experimental data (Zhang, Hu, Cheung and Chan, 2007; Hu and Hu, 2012; Bugni, Canay and Shaikh, 2018; Offer-Westort, Coppock and Green, 2021; Bai, Romano and Shaikh, 2022; Robertson, Lee, López-Kolkovska and Villar, 2023; Zhao, 2023; Wei, Ma and Wang, 2025; Bibaut and Kallus, 2025).
@@
Nevertheless, because CARA designs adjust treatment assignments based on observed participant responses, delays in outcome observation can significantly hinder their effectiveness. Intuitively, if primary outcomes for some participants remain unobserved at the time of interim analysis, the experimental designer may lack information needed to optimize treatment assignments for future participants with similar covariate profiles. These delays often depend on both the treatment arm and participant covariates, as is common in sequentially enrolled randomized experiments.
As an example, we revisit the study by Fahey, Njau, Katabaro, Mfaume, Ulenga, Mwenda et al. (2020), a randomized experiment conducted in the Shinyanga region of Tanzania between April 24 and December 14, 2018. The study aimed to assess the impact of cash incentives on retention in care and viral suppression among people living with HIV (PLHIV). A total of 530 PLHIV aged 18 or older were sequentially enrolled and randomly assigned to either a treatment group receiving cash incentives or a control group without any incentive. Viral load was measured six months after treatment assignment through blood draws. However, delays occurred, as the lab test required a separate clinical visit and some participants missed their scheduled appointments. We calculated the number of days delayed as the difference between the date of the viral load test and the six-month follow-up date, separately for two covariate strata defined by biological sex. From Figure 1, it is clear that the distribution of outcome delays varies significantly depending on treatment assignment and gender. Specifically, the treatment arm exhibits less delay, and within the treatment arm, the delayed response issue seems less pronounced among female participants. Overall, the empirical observation highlights the importance of accounting for delayed outcomes that are both arm- and covariate-specific.
Figure 1: Delays in viral load blood test across treatment arms and covariate strata (biological sex).
To our knowledge, although there has been prior work on developing valid statistical inference procedures in the presence of delayed responses in CARA (Bai, Hu and Rosenberger, 2002; Hu, Zhang, Cheung and Chan, 2008; Robertson, Lee, López-Kolkovska and Villar, 2023), no CARA designs has incorporated the delay mechanism when optimizing experimental objectives. In this paper, we address the challenges that response delays pose to covariate-adjusted response-adaptive (CARA) designs. We begin by examining whether existing optimal allocation strategies remain valid under delayed outcomes and, when they do not, derive the optimal allocation in such settings. We then develop methods to adapt CARA designs accordingly. Finally, we propose an efficient estimator that accounts for delayed responses at the end of the experiment. We elaborate on these contributions in more detail below.
@@
First, we demonstrate that the classical optimal allocation strategy used in existing CARA designs no longer provides optimal treatment assignment when accounting for arm- and covariate-dependent delays. Specifically, we first derive a semiparametric efficiency bound for treatment effect estimation in the presence of outcome delay (Theorem 1), which differ substantially from existing results in the literature (Hahn, 1998; Hirano, Imbens and Ridder, 2003; Cattaneo, 2010; Zhu and Zhu, 2023; Wu, Zheng, Zhang, Zhang and Wang, 2025). See Van der Vaart (2000) and Tsiatis (2006) for general discussions on semiparametric efficiency. This result allows us to show that the optimal delay mechanism involves a unique trajectory of treatment assignment probabilities that vary across different stages, resulting in treatment allocations with lower variance for power maximization, and enhanced failure reduction in welfare improvement designs. To our knowledge, this is the first design to explicitly explore delay mechanisms to optimize trial objectives, while existing literature focuses on conducting sequential tests (Hampson and Jennison, 2013), designing Bayesian adaptive designs where delay is independent of the arm and covariates (Lin, Thall and Yuan, 2020), or conducting statistical inference after experiments have concluded (Bai, Hu and Rosenberger, 2002; Hu, Zhang, Cheung and Chan, 2008).
@@
Second, since the delay mechanism is unknown a priori and remains only partially estimable during the experiment, we propose a fully forward-looking CARA design that sequentially updates treatment assignment probabilities to target a trial objective at the conclusion of the experiment. In this process, our algorithm offers several extrapolation strategies when estimating delay mechanism, allowing better alignment with the designer’s prior knowledge. We further discuss in detail two design objectives frequently adopted in current CARA literature: maximizing experiment power (Tymofyeyev, Rosenberger and Hu, 2007; Zhao, 2023) and improving overall participant welfare (or failure reduction) while maintaining statistical power constraints (Rosenberger, Stallard, Ivanova, Harper and Ricks, 2001; Robertson, Lee, López-Kolkovska and Villar, 2023). We justify the benefits of our proposed forward-looking CARA design through theoretical investigation (Section 4) and a synthetic case study (Section 5). Under the design objective of maximizing experiment power, our proposed forward-looking CARA design attains higher estimation efficiency compared to the Neyman allocation and complete randomization designs in the presence of delayed responses. Under the design objective of improving overall participant welfare, we demonstrate that our proposed forward-looking CARA design achieves additional failure reduction, leading to greater welfare improvement compared to both the ethical design and the complete randomization design while maintaining statistical power. For both design objectives, our proposed design can provide valid statistical inference, which is justified in Theorem 3.
@@
Third, we make theoretical contributions to the literature on CARA and adaptive experimental designs by establishing general conditions under which estimated treatment effects are consistent and asymptotically normally distributed (Theorem 2). Specifically, we do not restrict the potential outcomes or the delay mechanism to any parametric family of distributions, thereby alleviating the burden of choosing a particular set of parametric assumptions. Our theoretical development builds on martingale methods (Hall and Heyde, 1980) and a high-level condition (Condition 1) that the optimized allocations converge in large samples, an assumption satisfied by many adaptive experimental designs. As a result, the general consistency and asymptotic normality findings may be of independent interest. We then specialize in our proposed design and verify the design consistency condition (Section 4.3), providing a suite of tools for designing and analyzing adaptive experiments with delayed outcomes, along with rigorous statistical guarantees.
1.2 Setup: CARA with arm- and covariate-dependent delays
@@
In the remainder of this introduction, we lay out the setup of CARA when responses are subject to delay. Formal statements of our assumptions are collected in Section 4. The experiment is conducted over $T$ stages, labeled $t=1,2,\dots,T$. In each stage $t$, $n_{t}$ participants are enrolled. Participants are indexed by $i=1,2,\dots,N$, with $N=\sum_{t=1}^{T}n_{t}$ being the total sample size. We use $N_{t}$ to denote the cumulative sample size at the end of stage $t$, that is, $N_{t}=\sum_{s=1}^{t}n_{s}$. Therefore, $N=N_{T}$, where we omit the subscript to save notation. Finally, $S_{i}=t$ denotes that individual $i$ is enrolled in stage $t$.
@@
Upon enrollment, baseline covariates $X_{i}\in\mathcal{X}$ are collected for each participant, and they are then randomized into one of the treatment arms, denoted by $A_{i}\in\mathcal{A}=\{0,1\}$, which represents their actual treatment status. Following the Neyman-Rubin causal model, the potential outcomes are denoted by $Y_{i}(0)$ and $Y_{i}(1)$. In line with existing literature in adaptive design, we are in a scenario where the potential outcomes and the covariates, $(Y_{i}(0),Y_{i}(1),X_{i})$, are independent and identically distributed across different stages. For future reference, we also introduce notation for mean potential outcomes: $\mu(x,a)=\mathbb{E}[Y_{i}(a)|X_{i}=x]$ and $\mu(a)=\mathbb{E}[Y_{i}(a)]$, and the conditional and unconditional treatment effects are defined as $\tau(x)=\mu(x,1)-\mu(x,0)$ and $\tau=\mu(1)-\mu(0)$. Other moments of the potential outcomes are $\mu_{s}(x,a)=\mathbb{E}[Y_{i}(a)^{s}|X_{i}=x]$ and $\mu_{s}(a)=\mathbb{E}[Y_{i}(a)^{s}]$, but we omit the subscript whenever $s=1$ for ease of notation.
@@
In CARA designs, the treatment assignment is sequentially updated based on accumulated data to achieve a pre-specified design objective. Specifically, at stage $t$, suppose we have collected historical participant information, denoted as $\mathcal{H}_{t}=\{(A_{i},X_{i},Y_{i},D_{i})\}_{i=1}^{N_{t}}$, where we recall that the sample indexed by $i=1,2,\dots,N_{t}$ consists of all participants enrolled in the first $t$ stages. The treatment assignment for the next stage, $t+1$, denoted as $\hat{e}^{*}_{t+1}(a|x)$, is determined by both this historical information and the covariate information of the newly enrolled participants in Stage $t+1$. Formally, this is captured by $\mathbb{P}[A_{i}=a|X_{i}=x,S_{i}=t+1,\mathcal{H}_{t}]=\hat{e}^{*}_{t+1}(a|x)$, which means the treatment assignment adapts to $\mathcal{H}_{t}$.
In this manuscript, we work under the setting that $T$ and $n_{t}$ are pre-fixed. We also hope to note that, in general adaptive trial designs, determining the appropriate sample size and the number of stages involves balancing statistical rigor, operational feasibility, and ethical considerations. Before an experiment starts, practitioners often pre-specify the desired power and significance level, along with the expected treatment effect size based on prior studies or clinical knowledge, to compute an initial overall sample size. Then, the number of stages is selected by evaluating logistical considerations (e.g., recruitment rate, data availability, cost constraints). It is often the case that fewer stages are used to simplify logistics and reduce operational complexity, while more stages increase flexibility but impose higher management costs. Once the number of stages is decided, sample size allocation per stage can be determined either equally or unequally (e.g., adaptive design with small initial pilot study) based on anticipated interim analyses. It is also common practice to conduct simulations during the planning stage to evaluate various stage/sample size configurations, ensuring feasibility and ethical appropriateness of the chosen adaptive strategy. Our design thus aligns with the case where there is a finite number of stages and each stage contains an equal/unequal number of participants.
@@ 2 Optimal delay-adjusted treatment allocation in CARA
While the design goals of CARA vary across applications, maintaining sufficient statistical power to assess the treatment’s effectiveness on the primary outcome remains a central concern. A statistically efficient estimator of the treatment effect also supports a robust adaptive design, as the estimated effects guide the sequential revision of the treatment allocation strategy. However, it remains unclear what constitutes a good estimator in the presence of delayed outcomes. Moreover, the semiparametric efficiency bound for estimating the treatment effect in multi-stage experiments with delayed responses has not yet been established. To provide practical guidance for designing response-adaptive experiments, we begin by presenting a new result on the semiparametric efficiency bound for estimating the average treatment effect:
$\int\limits_{x\in\mathcal{X}}p(x)\Bigg{[}\frac{\sigma^{2}(x,1)}{\sum\limits_{t=1}^{T}r_{t}\rho(T-t|x,1)e_{t}(1|x)}$ $+\frac{\sigma^{2}(x,0)}{\sum\limits_{t=1}^{T}r_{t}\rho(T-t|x,0)e_{t}(0|x)}+\big{(}\tau(x)-\tau\big{)}^{2}\Bigg{]}\mbox{d}x,$ (1)
@@
where $r_{t}=\mathbb{P}[S_{i}=t]$ being the fraction of participants enrolled in stage $t$, $p(x)$ is the probability density function of the covariate $X$, and $\sigma^{2}(x,a)$ is the conditional variance of the potential outcome $Y_{i}(a)$. See Theorem 1 in Section 4 below for a rigorous statement.
To intuitively understand Eq (1), it helps to first revisit the classical semiparametric efficiency bound in a static (one-period) setting without delays, which is given by:
$\int\limits_{x\in\mathcal{X}}p(x)\Bigg{[}\frac{\sigma^{2}(x,1)}{e(1|x)}$ $+\frac{\sigma^{2}(x,0)}{e(0|x)}+\big{(}\tau(x)-\tau\big{)}^{2}\Bigg{]}\mbox{d}x.$ (2)
In this classical formulation, the denominators (i.e, propensity scores $e_{t}(a|x)$) adjust for covariate-specific treatment assignment. The key difference between Eq (2) and Eq (1) is the introduction of an additional term $\rho(T-t|x,a)$ in the denominators, which accounts for the probability distribution of response delays. Intuitively, delays introduce extra uncertainty or missingness in observing outcomes. As we operate in a multiple-stage experimental setting, we aggregate these delay-related probabilities across all stages, appropriately weighted by each stage’s sample size captured by $r_{t}$. Hence, the revised efficiency bound appropriately adjusts for both the random assignment of treatments (captured by propensity scores) and the randomness arising from delays in observing responses.
@@
From the above result, it is evident that when the delay mechanism of the response depends on the covariates and treatment status, the optimal treatment allocation rules in classical CARA designs must be modified to account for outcome delays. In what follows, we examine how the optimal treatment allocation rule is affected by the arm- and/or covariate-dependent delayed responses in two common experimental objectives: one aimed at enhancing statistical power (Tymofyeyev, Rosenberger and Hu, 2007), and the other focused on minimizing the expected number of failures while maintaining a minimum power requirement for binary outcomes (Rosenberger, Stallard, Ivanova, Harper and Ricks, 2001). In both design goals, the variance of the estimated treatment effects plays an essential role. To simplify the presentation, we focus on two-arm CARA designs with $a\in\{0,1\}$ in this paper. For multi-arm experiments, the design objective can be generalized using non-centrality parameters, as discussed in Lachin (1977) and Tymofyeyev, Rosenberger and Hu (2007).
@@
Design objective 1: Optimal treatment allocation for power maximization. The objective is to sequentially assign treatments to minimize the variance of the estimated average treatment effect (ATE), thereby maximizing statistical power. The optimal allocation is the solution to the optimization problem:
@@
$\min_{e_{1}(\cdot),\dots,e_{T}(\cdot)}\ \mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))\quad\text{subject to}\quad\delta\leq e_{t}(\cdot)\leq 1-\delta\quad\text{for }t=1,\ldots,T,$ (3)
@@
where $\delta$ ensures that the assignment probabilities are bounded away from 0 and 1, maintaining sufficient randomness in treatment assignments (Ma and Wang, 2020; Heiler and Kazak, 2021; Sasaki and Ura, 2022; Ma, Sasaki and Wang, 2025; Dorn, 2025). The objective function $\mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))$ is a component of the semiparametric efficiency bound related to the propensity scores, and is defined as:
@@
$\mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))=\sum_{x\in\mathcal{X}}p(x)\left[\frac{\sigma_{t}^{2}(x,1)}{\sum_{t=1}^{T}r_{t}\rho(T-t|x,1)e_{t}(x)}+\frac{\sigma_{t}^{2}(x,0)}{\sum_{t=1}^{T}r_{t}\rho(T-t|x,0)(1-e_{t}(x))}\right].$